OpenStudy (ayyookyndall):

**WILL MEDAL AND FAN!!! What is an equation for the line that passes through points (-1,-4) and (1,4)? Write an equation in slope-intercept form. ** Please explain

OpenStudy (ayyookyndall):

@aum @Walkerman15 @satellite73 @Destinymasha @bibby

OpenStudy (nikato):

do you know how to find the slope of the line that passes through those 2 points?

OpenStudy (ayyookyndall):

No. @nikato

OpenStudy (nikato):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }= slope\]

OpenStudy (nikato):

and you're given two points (x1,y1) and (x2,y2) plug in your points and solve for slope

OpenStudy (ayyookyndall):

So -4 - 4 = -8 And -1 - 1 = -2 @nikato

OpenStudy (nikato):

Yes. And what's -8/-2 ?

OpenStudy (ayyookyndall):

-4. So how would I write that problem out?

OpenStudy (nikato):

It's actually 4. Negative number divided by negative=positive So now you found that the slope is 4. Do you know what the slope intercept form is?

OpenStudy (ayyookyndall):

No.

OpenStudy (nikato):

Well it looks like this y=mx+b Where m= slope b= y-intercept And you have a point on the line (x,y)

OpenStudy (ayyookyndall):

y= 4x-2?

OpenStudy (nikato):

And you know that m=4 And you could use either one of the points for (x,y) Plug the numbers in and try to solve for b

OpenStudy (ayyookyndall):

B would be -2

OpenStudy (nikato):

No. B is not -2.

OpenStudy (ayyookyndall):

Wait I did this all wrong y2 is 4 and y1 is -4 it would be 8

OpenStudy (nikato):

Lets use (1,4) So plug in to y=mx+b 4= (4)(1) +b 4=4+b -4. -4 ---------- 0=b

OpenStudy (nikato):

Or if you used the other -4= (4)(-1)+b -4=-4+b +4. +4 ----------- 0=b

OpenStudy (ayyookyndall):

I did it wrong. It should be positive 8 and positive 2 and that equals 4

OpenStudy (nikato):

What? I've already showed you. b=0 ^^^^^

OpenStudy (ayyookyndall):

The y's equals 8 so that would be the y intercept.

OpenStudy (ayyookyndall):

So the answer is .....y = 4x+0

OpenStudy (nikato):

Yes. That is correct. But you could omit the +0 And just write y=4x, it's the same thing

OpenStudy (ayyookyndall):

How did you get 0 though?

OpenStudy (nikato):

Did you see above? I've showed you step by step by plugging into either points but you only had to do one

OpenStudy (ayyookyndall):

Oh yeah. Can you help me with one more?

OpenStudy (ayyookyndall):

@nikato

OpenStudy (nikato):

So you get it now? And yea, I can help you

OpenStudy (ayyookyndall):

Yes. Thank you. Which lines are parallel if m<4 = m<5? Justify your answer. Please Explain** https://www.connexus.com/content/media/336756-242011-55159-PM-1555594245.gif

OpenStudy (nikato):

Can you screen shot or something It says I need to log in

OpenStudy (nikato):

Ok. What do you think?

OpenStudy (ayyookyndall):

Is it referring to the numbers or letters?

OpenStudy (ayyookyndall):

I don't get it 4 and 5 are on 2 different lines.

OpenStudy (ayyookyndall):

@nikato do you understand?

OpenStudy (nikato):

Sorry. Was watching the baseball game. Giants just won! And the letters refer to the lines, but the numbers refer to the angle

OpenStudy (nikato):

Have you heard of the alternate interior theorem

OpenStudy (ayyookyndall):

Yes, but I don't know what it does.

OpenStudy (ayyookyndall):

So it would parallel to r and s?

OpenStudy (nikato):

Yes. That is right. r and s would be parallel

OpenStudy (ayyookyndall):

R and S would be parallel because of the alternate interior theorem? @nikato

OpenStudy (nikato):

Correct

OpenStudy (ayyookyndall):

Okay one more question please. @nikato

OpenStudy (ayyookyndall):

Find the values of x, y, z. The diagram is not to scale. I already found x. I need to find y and z. How?

OpenStudy (nikato):

Sorry was AFK. So after you find x. Solve 180-x to find z because x and z are supplementary which means they add up to 180. And once you find z, remember all 3 angles of a triangle add up to 180. This should help you find y

OpenStudy (ayyookyndall):

z is 99? Right? @nikato

OpenStudy (nikato):

Correct

OpenStudy (ayyookyndall):

So z is 68?

OpenStudy (ayyookyndall):

Wow thank you so much!! You a lifesaver!!!! @nikato

OpenStudy (nikato):

Correct. And no problem @ayyookyndall Happy to help you. You're a quick learner